# A Survey of Reinforcement Learning For Economics

### 23 Mar 2026 v4

@@ 2.2 Overlapping Terminology
The term contextual bandit is especially liable to misreading by economists. To an economist, the “context” is simply the state variable $x_{t}$, and a contextual bandit looks like an MDP with an unknown reward parameter. What the RL literature signals by “context” is a specific structural restriction on transitions, the agent’s action has no causal effect on the next context, so that $P(x_{t+1}\mid x_{t},a_{t})=P(x_{t+1})$. This exogeneity assumption separates the exploration problem (learning which arm is best given the current context) from the planning problem (choosing actions that influence future states). When contexts evolve exogenously, there is no long-horizon credit assignment, and the problem reduces to repeated one-period optimization under uncertainty. The term “contextual” therefore flags a modeling assumption about dynamics, not merely the presence of observable covariates.
@@
A concrete example clarifies how the two culture see a common problem. The RL literature frames some recommender systems as contextual bandits, where user covariates are the context, the recommendation is the arm, and a click or rating is the reward. An economist might instead view movie recommendation as a two-sided learning problem in which the platform learns user preferences while users simultaneously explore the catalog and update their own tastes. The bandit formulation absorbs the user’s utility maximization into the environment’s reward signal and treats user arrivals as exogenous. This is a modeling choice, not a fact about the world. Also, the object called a “bandit” in this formulation, a one-step decision under exogenous context, is a slightly different mathematical object than the Bayesian sequential allocation problem of Gittins (1979), even though both carry the same name and are related.  [^6] Lattimore (2016) proves that the Gittins index with a flat Gaussian prior achieves finite-time regret of the same order as UCB, and that the index decomposes as posterior mean plus an exploration bonus that shrinks toward zero near the horizon, structurally resembling but differing from the UCB confidence bound. As apparent, the RL view often abstracts away from a lot of complexity in the world (human learning, human interaction) to squash the problem into the straightforward language of optimal control ($\pi$, $V$, $P,r$), while the economics literature prefers to allow for much more complex models of human-firm interaction.
@@
The word policy is itself overloaded, and the divergence runs deeper than terminology. In economics, a “policy” is a rule set by a government, central bank, or regulator, a tax schedule, a subsidy, an interest rate rule, a licensing requirement. These rules are part of the environment in which private agents (consumers, firms) optimize. “Policy evaluation” in this tradition means changing some aspect of the environment and asking how agents would respond: if the earned income tax credit were expanded, how would labor supply shift? The standard workflow proceeds in two steps: first estimate or calibrate the structural model (preferences, technology, market interactions), then simulate counterfactuals under the alternative rule. In RL, “policy” means the agent’s own decision rule $\pi(a|s)$, and “policy evaluation” means computing $V^{\pi}(s)$, the expected return from following that rule in a fixed environment. RL typically assumes access to a high-fidelity simulator, a physics engine, a game, a digital twin, whose rules do not change; the interesting question is how the agent should behave within those rules, not what happens if the rules change. When the environment does shift, RL treats it as a nuisance (sim-to-real transfer, domain adaptation) rather than the object of study; in offline RL (Section 11), distributional shift between the training data and the deployed environment becomes a central concern, bringing RL closer to the economist’s counterfactual reasoning. [^7] A notable exception is Tomasev et al. (2020), where AlphaZero was used to evaluate alternative chess rule sets proposed by Kramnik, asking how optimal play changes under modified game rules. This is precisely an environment counterfactual of the kind economists routinely study.
@@
Identification carries fundamentally different weight across the two fields. In econometrics, a parameter $\theta$ is identified if it is uniquely pinned down by the combination of the data-generating process and the model’s maintained assumptions; formally, no two distinct parameter values $\theta\neq\tilde{\theta}$ can generate the same distribution of observable data (Lewbel, 2019). Identification is a logical property of the model, not a statistical one; if identification fails, no amount of additional data or more sophisticated estimation will recover the parameter, because multiple parameter values are observationally equivalent. In RL, there is little general identification discourse. The closest analogs are narrow and domain-specific. In inverse reinforcement learning, reward identifiability asks whether the reward function can be uniquely recovered from observed optimal behavior; Kim et al. (2021) formalize this for MaxEnt MDP models, showing that for deterministic dynamics, strong identifiability requires the domain graph to be coverable and aperiodic. [^8] Kim et al. (2021) explicitly note that the econometrics literature on identifying dynamic discrete choice models (Rust, 1994) addresses “an equivalent problem” to reward identifiability in IRL, providing a direct bridge between the two fields. In model-based RL, system identification refers to learning the transition dynamics $\hat{T}$ well enough for the resulting policy to perform well, a usage inherited from control theory rather than econometrics (Ross and Bagnell, 2012). RL researchers focus on out-of-sample performance of the learned controller, so whether parameters are identified in the econometric sense matters less; what matters is that the policy generalizes. [^9] The one RL subfield where identification is central, inverse reinforcement learning, is the subject of the companion survey (Rust and Rawat, 2026).
@@
The term regret itself diverges across fields. In economics, regret is either an emotion that modifies preferences under uncertainty (Loomes and Sugden 1982) or a static minimax decision criterion for choosing among actions when probabilities are unknown (Savage 1951, Manski 2004). In RL and computer science, regret is a dynamic performance metric, the cumulative gap between the agent’s returns and those of the best fixed policy in hindsight, and sublinear growth of this quantity is the standard benchmark for online learning algorithms. In economics, efficiency concerns welfare: Pareto efficiency asks whether any reallocation could improve one agent’s utility without harming another, and allocative efficiency asks whether goods flow to their highest-valued uses. In RL, efficiency concerns resources: sample efficiency measures how many environment interactions an algorithm needs to find a good policy, and computational efficiency measures the operations required per timestep.
@@
In economics, the discount factor is a structural parameter encoding time preference, an agent’s intrinsic willingness to trade present for future consumption. Koopmans (1960) derived it axiomatically from preference postulates, principally stationarity and impatience, showing that these behavioral axioms uniquely imply an exponential discounted utility representation with a single discount factor strictly between zero and one (Bleichrodt et al., 2008). Economists treat this parameter as a deep primitive of preferences, to be estimated from data on intertemporal choices, and deviations from constant discounting (such as the quasi-hyperbolic present bias of Laibson 1997) are studied as substantive behavioral phenomena. In RL, the discount factor has historically served a more pragmatic role, ensuring that infinite-horizon returns remain finite and that the Bellman operator is a contraction, thereby guaranteeing convergence of dynamic programming algorithms. It is typically treated as a hyperparameter to be tuned for computational performance rather than a structural claim about the agent’s preferences. Pitis (2019) bridges this gap, axiomatically deriving discounting in RL from rationality postulates and showing that the fixed discount factor is best understood as an optimizing representation rather than a literal description of time preference.
@@ 2.3 Structural Equivalences
Beyond terminological differences, several formal objects in RL and economics are mathematically identical. The softmax (or Boltzmann) policy used throughout RL is the multinomial logit model of McFadden (1974). The RL softmax policy selects actions according to
@@
The advantage function $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$ measures how much better action $a$ is compared to the average action under $\pi$. In the discrete choice framework, this is precisely the choice-specific value function net of the ex-ante value function, a quantity that appears in the Hotz and Miller (1993) CCP estimator for dynamic discrete choice models.
These equivalences reflect a common convex-analytic structure. The log-sum-exp value function and the negative Shannon entropy are Fenchel conjugates of each other, so maximizing expected payoffs with an entropy bonus and recovering utilities from observed choice probabilities are two views of the same optimization. Chiong et al. (2016) build on this conjugate duality to develop identification and estimation methods for dynamic discrete choice models, and Fosgerau and Sørensen (2022) show that the underlying structure extends well beyond the logit case.
The two fields arrived at this shared mathematics from opposite directions. In RL, entropy regularization began as a pragmatic trick to prevent premature convergence and encourage exploration (Williams and Peng, 1991), and only decades later did Ziebart (2010) and Levine (2018) provide decision-theoretic foundations showing that the resulting policies are robust to model misspecification. In economics, the softmax emerged not from any concern about exploration but from the random utility framework of McFadden (1974), where agents are fully informed and choose deterministically; the apparent randomness arises entirely from the econometrician’s inability to observe all relevant taste variation. The RL agent randomizes because it is ignorant of the environment; the economic agent appears to randomize because the observer is ignorant of the agent’s preferences.
@@ 2.4 Notation
Table 2: Notation mapping between reinforcement learning and economics.
RL Term Symbol Economics Equivalent Symbol
@@
Action $a\in\mathcal{A}$ Choice, control variable $d_{t}$, $u_{t}$ [^10] The letter $u$ appears twice in the economics column with different meanings: $u_{t}$ denotes the control variable (action), while $u(x,d)$ denotes per-period utility (reward). Context usually disambiguates, but readers should note the collision.
@@
Discount factor $\gamma$ [^11] RL permits $\gamma=0$ (myopic, one-step) and $\gamma=1$ (undiscounted, episodic tasks with guaranteed termination). Economics requires $\beta\in(0,1)$ strictly; the contraction mapping argument for the Bellman operator relies on $\beta<1$. Discount factor $\beta$
@@ 4.1.4 Williams (1992)
Consider a robot arm that must apply a continuous torque $a\in\mathbb{R}$ to reach a target angle. Q-learning requires computing $\max_{a}Q(s,a)$ at every update, which becomes a nested optimization problem  [^25] In continuous action spaces, $\max_{a}Q(s,a)$ has no closed-form solution in general and must be solved numerically at every Bellman update. Discretizing a $d$-dimensional action space on a grid of $m$ points per dimension costs $O(m^{d})$ evaluations per update step. when the action space is continuous. A policy gradient method sidesteps the issue. Parameterize the policy as a Gaussian $\pi_{\theta}(a|s)=\mathcal{N}(\mu_{\theta}(s),\,\sigma_{\theta}^{2}(s))$, [^26] The Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ has density $(2\pi\sigma^{2})^{-1/2}\exp(-(a-\mu)^{2}/2\sigma^{2})$. Here $\mu_{\theta}(s)$ and $\sigma_{\theta}(s)$ are neural network outputs parameterizing the policy mean and standard deviation. sample an action, observe the return, and update $\theta$ by REINFORCE. Virtually all continuous-control results in reinforcement learning descend from the policy gradient framework for this reason.
@@ 4.1.10 Fitted Value Iteration and Fitted Q-Iteration (2005)
Tabular Q-learning maintains a separate entry for every state-action pair. When $|\mathcal{S}|$ is large or the state space is continuous, as in most economic applications, this is infeasible. Fitted Q-Iteration (FQI) (Ernst et al., 2005) replaces the tabular update with a supervised regression step: given a batch of transitions, fit a function approximator to the Bellman targets.
@@ 4.2.1 Deep Q-Networks (2015)
DQN exceeded human-level performance on 29 of 49 games using a single architecture and hyperparameters. [^40] Hyperparameters are design choices fixed before training begins, such as network depth, learning rate, and replay buffer size; they are not estimated by gradient descent and are analogous to tuning parameters in nonparametric estimation (bandwidth, penalty weights). Games requiring long-horizon planning or sparse rewards, such as Montezuma’s Revenge, remained difficult.
@@ 5.5.1 Actor-Critic Architecture and Two-Timescale Convergence
The theoretical foundation is two-timescale stochastic approximation (Konda and Tsitsiklis, 2000), building on the two-timescale ODE convergence theory of Borkar (1997). [^98] The original analysis of Konda and Tsitsiklis (2000) uses the average-cost formulation with TD error $\delta_{t}=c(X_{t},U_{t})-\Lambda+V(X_{t+1})-V(X_{t})$, where $\Lambda$ is the average cost. The discounted variant presented here follows by replacing the average-cost baseline with $\gamma V(s^{\prime})$. A related but distinct ODE stability framework for single-timescale stochastic approximation, including Q-learning and TD learning, appears in Borkar and Meyn (2000). Run two concurrent learning processes:
@@
Under this separation, the actor sees a quasi-stationary critic: from the actor’s perspective, the critic provides approximately correct value estimates at each step. [^99] The two-timescale structure is analogous to nested optimization in structural estimation, where an inner loop solves for equilibrium given parameters and an outer loop searches over parameters. The critic’s inner loop (policy evaluation) must converge before the actor’s outer loop (policy improvement) takes a step. Wu et al. (2020) provide finite-time convergence rates ($\tilde{O}(\epsilon^{-2.5})$ sample complexity) for two-timescale actor-critic with linear approximation, and Tian et al. (2023) establish analogous rates for single-timescale actor-critic with multi-layer neural networks. The actor’s updates are then approximately unbiased policy gradient steps. Konda and Tsitsiklis (2000) prove convergence to a stationary point of $J(\omega)$ (i.e., $\nabla J(\omega)\to 0$).
Convergence requires a structural condition on the critic. The critic’s feature vectors must span the actor’s score functions $\nabla_{\omega}\log\pi_{\omega}(a|s)$, so that the critic’s approximation error lies orthogonal to the policy gradient direction. Under this compatibility condition, the critic’s projection error does not bias the actor’s gradient estimates.100100100The compatible function approximation theorem first appears in Sutton et al. (2000) and is the key structural requirement in Konda and Tsitsiklis (2000). It constrains the critic architecture to be “compatible” with the actor parameterization, the same condition that makes the natural policy gradient equal to the critic’s weight vector in Kakade (2002).
A2C (Advantage Actor-Critic) is the synchronous variant: collect a batch of transitions, compute TD errors, and update both networks. A3C (Mnih et al., 2016) parallelizes this across multiple workers updating a shared parameter server asynchronously.101101101Parallel workers decorrelate gradient estimates by exploring different parts of the state space simultaneously, removing the need for an experience replay buffer. Rigorous convergence theory for A3C’s lock-free asynchronous parameter updates remains an open problem; existing analyses of asynchronous stochastic approximation (Qu and Wierman, 2020) address classical asynchrony (different state-action pairs updated at different times on a single trajectory), not the parallel-worker gradient setting.

### 24 Mar 2026 v5

@@ 2.2 Overlapping Terminology
The term contextual bandit is especially liable to misreading. To a reader from the inference tradition, the “context” is simply the state variable $x_{t}$, and a contextual bandit looks like an MDP with an unknown reward parameter. What the RL literature signals by “context” is a specific structural restriction on transitions, the agent’s action has no causal effect on the next context, so that $P(x_{t+1}\mid x_{t},a_{t})=P(x_{t+1})$. This exogeneity assumption separates the exploration problem (learning which arm is best given the current context) from the planning problem (choosing actions that influence future states). When contexts evolve exogenously, there is no long-horizon credit assignment, and the problem reduces to repeated one-period optimization under uncertainty. The term “contextual” therefore flags a modeling assumption about dynamics, not merely the presence of observable covariates.
@@
The RL literature frames some recommender systems as contextual bandits, where user covariates are the context, the recommendation is the arm, and a click or rating is the reward. One might instead view movie recommendation as a two-sided learning problem in which the platform learns user preferences while users simultaneously explore the catalog and update their own tastes. The bandit formulation absorbs the user’s utility maximization into the environment’s reward signal and treats user arrivals as exogenous. This is a modeling choice, not a fact about the world. Also, the object called a “bandit” in this formulation, a one-step decision under exogenous context, is a slightly different mathematical object than the Bayesian sequential allocation problem of Gittins (1979), even though both carry the same name and are related.  [^6] Lattimore (2016) proves that the Gittins index with a flat Gaussian prior achieves finite-time regret of the same order as UCB, and that the index decomposes as posterior mean plus an exploration bonus that shrinks toward zero near the horizon, structurally resembling but differing from the UCB confidence bound. RL abstracts away much of this complexity (human learning, strategic interaction) to fit the problem into the standard MDP framework ($\pi$, $V$, $P$, $r$).
@@
The word policy is overloaded. In the inference tradition, a “policy” is a rule set by a government, central bank, or regulator, a tax schedule, a subsidy, an interest rate rule, a licensing requirement. These rules are part of the environment in which private agents (consumers, firms) optimize. “Policy evaluation” in this tradition means changing some aspect of the environment and asking how agents would respond: if the earned income tax credit were expanded, how would labor supply shift? The standard workflow proceeds in two steps: first estimate or calibrate the structural model (preferences, technology, market interactions), then simulate counterfactuals under the alternative rule. In RL, “policy” means the agent’s own decision rule $\pi(a|s)$, and “policy evaluation” means computing $V^{\pi}(s)$, the expected return from following that rule in a fixed environment. RL typically assumes access to a high-fidelity simulator, a physics engine, a game, a digital twin, whose rules do not change; the interesting question is how the agent should behave within those rules, not what happens if the rules change. When the environment does shift, RL treats it as a nuisance (sim-to-real transfer, domain adaptation) rather than the object of study; in offline RL (Section 11), distributional shift between the training data and the deployed environment becomes a central concern, bringing RL closer to standard counterfactual reasoning. [^7] A notable exception is Tomasev et al. (2020), where AlphaZero was used to evaluate alternative chess rule sets proposed by Kramnik, asking how optimal play changes under modified game rules. This is precisely an environment counterfactual.
@@
Identification means different things in the two fields. In statistics, a parameter $\theta$ is identified if it is uniquely pinned down by the combination of the data-generating process and the model’s maintained assumptions; formally, no two distinct parameter values $\theta\neq\tilde{\theta}$ can generate the same distribution of observable data (Lewbel, 2019). Identification is a logical property of the model, not a statistical one; if identification fails, no amount of additional data or more sophisticated estimation will recover the parameter, because multiple parameter values are observationally equivalent. In RL, there is little general identification discourse. The closest analogs are narrow and domain-specific. In inverse reinforcement learning, reward identifiability asks whether the reward function can be uniquely recovered from observed optimal behavior; Kim et al. (2021) formalize this for MaxEnt MDP models, showing that for deterministic dynamics, strong identifiability requires the domain graph to be coverable and aperiodic. [^8] Kim et al. (2021) explicitly note that the literature on identifying dynamic discrete choice models (Rust, 1994) addresses “an equivalent problem” to reward identifiability in IRL, providing a direct bridge between the two fields. In model-based RL, system identification refers to learning the transition dynamics $\hat{T}$ well enough for the resulting policy to perform well, a usage inherited from control theory rather than statistics (Ross and Bagnell, 2012). RL researchers focus on out-of-sample performance of the learned controller, so whether parameters are identified in the statistical sense matters less; what matters is that the policy generalizes. [^9] The one RL subfield where identification is central, inverse reinforcement learning, is the subject of the companion survey (Rust and Rawat, 2026).
@@
Regret diverges across fields. In decision theory, regret is either an emotion that modifies preferences under uncertainty (Loomes and Sugden 1982) or a static minimax decision criterion for choosing among actions when probabilities are unknown (Savage 1951, Manski 2004). In RL and computer science, regret is a dynamic performance metric, the cumulative gap between the agent’s returns and those of the best fixed policy in hindsight, and sublinear growth of this quantity is the standard benchmark for online learning algorithms. In the other tradition, efficiency concerns welfare: Pareto efficiency asks whether any reallocation could improve one agent’s utility without harming another, and allocative efficiency asks whether goods flow to their highest-valued uses. In RL, efficiency concerns resources: sample efficiency measures how many environment interactions an algorithm needs to find a good policy, and computational efficiency measures the operations required per timestep.
@@
In the inference tradition, the discount factor is a structural parameter encoding time preference, an agent’s intrinsic willingness to trade present for future consumption. Koopmans (1960) derived it axiomatically from preference postulates, principally stationarity and impatience, showing that these behavioral axioms uniquely imply an exponential discounted utility representation with a single discount factor strictly between zero and one (Bleichrodt et al., 2008). This parameter is treated as a deep primitive of preferences, to be estimated from data on intertemporal choices, and deviations from constant discounting (such as the quasi-hyperbolic present bias of Laibson 1997) are studied as substantive behavioral phenomena. In RL, the discount factor has historically served a more pragmatic role, ensuring that infinite-horizon returns remain finite and that the Bellman operator is a contraction, thereby guaranteeing convergence of dynamic programming algorithms. It is typically treated as a hyperparameter to be tuned for computational performance rather than a structural claim about the agent’s preferences. Pitis (2019) bridges this gap, axiomatically deriving discounting in RL from rationality postulates and showing that the fixed discount factor is best understood as an optimizing representation rather than a literal description of time preference.
@@ 2.3 Structural Equivalences
Beyond terminological differences, several formal objects in RL and the inference tradition are mathematically identical. The softmax (or Boltzmann) policy used throughout RL is the multinomial logit model of McFadden (1974). The RL softmax policy selects actions according to
@@
The action-value function $Q^{\pi}(s,a)$ is the choice-specific value function $v_{\theta}(x,d)$ of the DDC literature (Table 2). The advantage function $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$ is therefore the choice-specific value net of the ex-ante value, a quantity that appears in the Hotz and Miller (1993) CCP estimator for dynamic discrete choice models. [^10] These equivalences reflect a common convex-analytic structure. The log-sum-exp value function and the negative Shannon entropy are Fenchel conjugates of each other, so maximizing expected payoffs with an entropy bonus and recovering utilities from observed choice probabilities are two views of the same optimization. Chiong et al. (2016) build on this conjugate duality to develop identification and estimation methods for dynamic discrete choice models, and Fosgerau and Sørensen (2022) show that the underlying structure extends well beyond the logit case.
The two fields arrived at this shared mathematics from opposite directions. In RL, entropy regularization began as a pragmatic trick to prevent premature convergence and encourage exploration (Williams and Peng, 1991), and only decades later did Ziebart (2010) and Levine (2018) provide decision-theoretic foundations showing that the resulting policies are robust to model misspecification. In the inference tradition, the softmax emerged not from any concern about exploration but from the random utility framework of McFadden (1974), where agents are fully informed and choose deterministically; the apparent randomness arises entirely from the econometrician’s inability to observe all relevant taste variation. The RL agent randomizes because it is ignorant of the environment; the economic agent appears to randomize because the observer is ignorant of the agent’s preferences.
@@ 2.4 Notation
Table 2: Notation mapping between reinforcement learning and the inference tradition.
RL Term Symbol Inference Tradition Symbol
@@
Action $a\in\mathcal{A}$ Choice, control variable $d_{t}$, $u_{t}$ [^11] The letter $u$ appears twice in the inference-tradition column with different meanings: $u_{t}$ denotes the control variable (action), while $u(x,d)$ denotes per-period utility (reward). Context usually disambiguates, but readers should note the collision.
@@
Discount factor $\gamma$ [^12] RL permits $\gamma=0$ (myopic, one-step) and $\gamma=1$ (undiscounted, episodic tasks with guaranteed termination). The inference tradition requires $\beta\in(0,1)$ strictly; the contraction mapping argument for the Bellman operator relies on $\beta<1$. Discount factor $\beta$
@@ 4.1.4 Williams (1992)
Consider a robot arm that must apply a continuous torque $a\in\mathbb{R}$ to reach a target angle. Q-learning requires computing $\max_{a}Q(s,a)$ at every update, which becomes a nested optimization problem  [^26] In continuous action spaces, $\max_{a}Q(s,a)$ has no closed-form solution in general and must be solved numerically at every Bellman update. Discretizing a $d$-dimensional action space on a grid of $m$ points per dimension costs $O(m^{d})$ evaluations per update step. when the action space is continuous. A policy gradient method sidesteps the issue. Parameterize the policy as a Gaussian $\pi_{\theta}(a|s)=\mathcal{N}(\mu_{\theta}(s),\,\sigma_{\theta}^{2}(s))$, [^27] The Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ has density $(2\pi\sigma^{2})^{2}\exp(-(a-\mu)^{2}/2\sigma^{2})$. Here $\mu_{\theta}(s)$ and $\sigma_{\theta}(s)$ are neural network outputs parameterizing the policy mean and standard deviation. sample an action, observe the return, and update $\theta$ by REINFORCE. Virtually all continuous-control results in reinforcement learning descend from the policy gradient framework for this reason.
@@ 4.1.10 Fitted Value Iteration and Fitted Q-Iteration (2005)
Tabular Q-learning maintains a separate entry for every state-action pair. When $|\mathcal{S}|$ is large or the state space is continuous, as in most applied problems, this is infeasible. Fitted Q-Iteration (FQI) (Ernst et al., 2005) replaces the tabular update with a supervised regression step: given a batch of transitions, fit a function approximator to the Bellman targets.
@@ 4.2.1 Deep Q-Networks (2015)
DQN exceeded human-level performance on 29 of 49 games using a single architecture and hyperparameters. [^41] Hyperparameters are design choices fixed before training begins, such as network depth, learning rate, and replay buffer size; they are not estimated by gradient descent. Games requiring long-horizon planning or sparse rewards, such as Montezuma’s Revenge, remained difficult.
@@ 4.2.5 AlphaGo Zero (2017)
4.2.6 Decision Transformers (2021)
Chen et al. (2021) proposed replacing Bellman backups with autoregressive sequence modeling. The Decision Transformer conditions a causal GPT-style Transformer on trajectories $\tau=(\hat{R}_{1},s_{1},a_{1},\hat{R}_{2},s_{2},a_{2},\ldots)$, where $\hat{R}_{t}$ is the return-to-go (desired future cumulative reward). At test time, conditioning on a high target return extracts a high-performing policy without any temporal-difference learning. Janner et al. (2021) extended this to the Trajectory Transformer, modeling entire trajectories as flat token sequences with continuous dimensions discretized into bins, enabling planning via beam search over trajectories.
The approach has fundamental limitations that Bellman-based methods do not share. Brandfonbrener et al. (2022) proved that return-conditioned supervised learning (RCSL) recovers optimal policies only under assumptions strictly stronger than those needed for dynamic programming: near-deterministic dynamics, expert data coverage, and a unique mapping from returns to optimal actions. In stochastic environments, high returns may reflect environmental luck rather than good decisions, causing RCSL to imitate lucky-but-suboptimal trajectories. Paster et al. (2022) demonstrated this concretely: in a simple gambling MDP, the Decision Transformer conditioned on high returns selects risky gambles over the optimal safe action, even with infinite data. [^56] Emmons et al. (2022) showed that a simple two-layer MLP matches the Transformer architecture on D4RL benchmarks, suggesting the autoregressive structure provides conditional density estimation rather than temporal reasoning. The essential element is conditioning on the right outcome variable, not the architecture.
A second limitation is that RCSL cannot stitch suboptimal trajectory segments. If the dataset contains two trajectories that each visit a useful intermediate state but from different starting points, Bellman-based methods can compose the better segments by propagating values backward through the shared state. Sequence models, which predict forward autoregressively, cannot perform this backward composition.
@@ 5.5.1 Actor-Critic Architecture and Two-Timescale Convergence
The theoretical foundation is two-timescale stochastic approximation (Konda and Tsitsiklis, 2000), building on the two-timescale ODE convergence theory of Borkar (1997).100100100The original analysis of Konda and Tsitsiklis (2000) uses the average-cost formulation with TD error $\delta_{t}=c(X_{t},U_{t})-\Lambda+V(X_{t+1})-V(X_{t})$, where $\Lambda$ is the average cost. The discounted variant presented here follows by replacing the average-cost baseline with $\gamma V(s^{\prime})$. A related but distinct ODE stability framework for single-timescale stochastic approximation, including Q-learning and TD learning, appears in Borkar and Meyn (2000). Run two concurrent learning processes:
@@
Under this separation, the actor sees a quasi-stationary critic: from the actor’s perspective, the critic provides approximately correct value estimates at each step.101101101The two-timescale structure is analogous to nested optimization in structural estimation, where an inner loop solves for equilibrium given parameters and an outer loop searches over parameters. The critic’s inner loop (policy evaluation) must converge before the actor’s outer loop (policy improvement) takes a step. Wu et al. (2020) provide finite-time convergence rates ($\tilde{O}(\epsilon^{-2.5})$ sample complexity) for two-timescale actor-critic with linear approximation, and Tian et al. (2023) establish analogous rates for single-timescale actor-critic with multi-layer neural networks. The actor’s updates are then approximately unbiased policy gradient steps. Konda and Tsitsiklis (2000) prove convergence to a stationary point of $J(\omega)$ (i.e., $\nabla J(\omega)\to 0$).
Convergence requires a structural condition on the critic. The critic’s feature vectors must span the actor’s score functions $\nabla_{\omega}\log\pi_{\omega}(a|s)$, so that the critic’s approximation error lies orthogonal to the policy gradient direction. Under this compatibility condition, the critic’s projection error does not bias the actor’s gradient estimates.102102102The compatible function approximation theorem first appears in Sutton et al. (2000) and is the key structural requirement in Konda and Tsitsiklis (2000). It constrains the critic architecture to be “compatible” with the actor parameterization, the same condition that makes the natural policy gradient equal to the critic’s weight vector in Kakade (2002).
A2C (Advantage Actor-Critic) is the synchronous variant: collect a batch of transitions, compute TD errors, and update both networks. A3C (Mnih et al., 2016) parallelizes this across multiple workers updating a shared parameter server asynchronously.103103103Parallel workers decorrelate gradient estimates by exploring different parts of the state space simultaneously, removing the need for an experience replay buffer. Rigorous convergence theory for A3C’s lock-free asynchronous parameter updates remains an open problem; existing analyses of asynchronous stochastic approximation (Qu and Wierman, 2020) address classical asynchrony (different state-action pairs updated at different times on a single trajectory), not the parallel-worker gradient setting.